# Alexandrov spaces are CS sets

### v1

@@ Abstract.
We prove that the extremal stratification of an Alexandrov space by Perelman–Petrunin is a CS stratification in the sense of Siebenmann. We also show that every space of directions of an Alexandrov space without proper extremal subsets is homeomorphic to a sphere. In the appendix we give an example of a primitive extremal subset of codimension $2$ that is not an Alexandrov space with respect to the induced intrinsic metric.
@@
Alexandrov spaces, stratification, CS sets, extremal subsets
@@ 1. Introduction
The extremal stratification of an Alexandrov space by Perelman–Petrunin is a CS stratification in the sense of Siebenmann.
@@
In fact, from the geometrical point of view, our proof is just a combination of known results and provides nothing new. Nevertheless, the author believes that it is worth publishing because it connects two concepts in different fields, yet even a statement cannot be found in the literature. Another motivation comes from the collapsing theory of Riemannian manifolds, where the extremal stratification of limit Alexandrov spaces plays an essential role ([28], [13], [14]; especially in the sheaf-theoretical conjectures by Alesker [1]).
@@
At the same time as [26], Perelman–Petrunin [29] introduced the notion of an extremal subset, which is a singular set of an Alexandrov space defined by a purely geometrical condition. By extending the Morse theory to extremal subsets, they proved that any Alexandrov space is, in a certain sense, uniquely stratified into its own extremal subsets. We refer to it as the extremal stratification in this paper. The extremal stratification is finer than the MCS stratification (the arrow (2)): the former reflects not only the topological singularity but also the geometrical singularity of an Alexandrov space. For example, the extremal stratification of a squire consists of the four vertices, the interiors of the four edges, and the interior of the squire itself, whereas the MCS stratification consists only of the boundary and the interior of the squire.
@@
The above-mentioned Morse theory relies heavily on the work of Siebenmann [34] on the deformation of homeomorphisms on stratified sets. In [34], Siebenmann introduced two notions of stratified sets: the one is a CS set (= cone-like stratified) and the other is a WCS set (= weakly cone-like stratified). As the names suggest, CS sets are WCS sets (the arrow (3)). From our viewpoint, the main difference of these two is whether the local normal trivialization preserves the stratification or not (see Remark 2.3). Still, the fundamental deformation theorem of Siebenmann works in the weaker setting of WCS sets.
@@
In [25], [26], Perelman only stated that the MCS stratification of an Alexandrov space is a WCS stratification (the arrow (4)), which was enough for him to develop the Morse theory by using the deformation theorem of Siebenmann for WCS sets. However, there was no mention of the CS structure in [25], [26], even in [29], and as already remarked, since then there have been no papers about it (except for a closely related result of Kapovitch [20, §9] mentioned below). This is why we consider this problem and the answer is, as stated in Theorem 1.1, given by the extremal stratification of Perelman–Petrunin (the arrow (1)).
@@
Actually the proof of Theorem 1.1 is almost the same as the original proof of the existence of the extremal stratification by Perelman–Petrunin [29, 3.8, 1.2]. The only difference is that we will make use of the relative fibration theorem with respect to extremal subsets developed later by Kapovitch [20, §9], instead of the absolute one given in [29, §2].
@@
The extremal stratification theory in particular implies that if an Alexandrov space contains no proper extremal subsets, then it must be a topological manifold. However, even if an Alexandrov space is a topological manifold, it does not necessarily mean that each link (i.e., the space of directions) is a topological sphere. See Example 2.22 based on the double suspension theorem of Cannon [7] and Edwards [9], [10]. The infinitesimal characterization of Alexandrov topological manifolds by Wu [35] only tells us that the link is a homology manifold with the homotopy type of a sphere (cf. [23]). Regarding the structure of the link, for now we will show the following simplest case (cf. [32, Ch. 3 Problem 32], [20]).
@@
Organization. In Section 2, we recall background material including the definitions of the four stratifications mentioned above. In Sections 3 and 4, we prove Theorems 1.1 and 1.2, respectively. In Appendix A, we give an example of a primitive extremal subset of codimension $2$ that is not an Alexandrov space with respect to the induced intrinsic metric, which also cannot be found in the literature.
Acknowledgments. I would like to thank Prof. Semyon Alesker for his initial question and further comments regarding the main theorem that motivated me to write this paper.
@@ 2. Preliminaries
Here we begin with the definition of a CS set and then recall some notions and results from Alexandrov geometry which will be used to prove the main theorem 1.1. For the latter, we will focus on stating their properties rather than giving the formal definitions. Actually, once those properties are assumed, the proof requires almost no geometrical knowledge.
@@ 2.1. CS sets
First we recall the definition of a CS set by Siebenmann [34, §1]. For simplicity, we restrict our attention to the finite-dimensional case.
@@
An isomorphism between (not necessarily TOP) stratified sets $X$ and $Y$ is a homeomorphism between $X$ and $Y$ preserving their strata of the same dimension, which will be denoted by $X\cong Y$.
For stratified sets $X$ and $Y$, the direct product $X\times Y$ has a natural stratification. The open cone $C(X):=X\times[0,1)/X\times\{0\}$ has a natural stratification as well, which is induced from the product structure of $X\times(0,1)$ together with the zero-dimensional stratum consisting only of the vertex of the cone. An open subset $U$ of $X$ also has a natural stratification by restriction.
@@
A TOP stratified set $X$ is called a CS set if for any $x\in X_{k}$ there exist an open neighborhood $U$ of $x$ in $X$ and a compact stratified set $L$ such that there is an isomorphism
@@
There is a weaker notion of a WCS set introduced by Siebenmann [34, §5]. Roughly speaking, there are two relaxations of the above conditions:
@@
Any Alexandrov space is known to be a WCS set, but not known to be a CS set in the literature. As we will see below, the problem lies in (2) rather than (1) (i.e., $C(L)$ is a real cone but $h$ might not be an isomorphism). In particular, we do not need the definition of the mock cone here.
@@ 2.2. Alexandrov spaces
An Alexandrov space is a complete finite-dimensional geodesic metric space with a lower curvature bound in terms of triangle comparison, introduced by Burago–Gromov–Perelman [6]. More precisely, any geodesic triangle is not thinner than the comparison triangle with the same sidelengths in the model plane of constant curvature. Typical examples are
  • Riemannian manifolds with sectional curvature bounded below;
  • their convex subsets and convex hypersurfaces;
  • their Gromov–Hausdorff limits;
  • their quotient spaces by isometric compact group actions.
See also [4], [3] for the formal definition and properties. Note that the dimension of an Alexandrov space means its Hausdorff dimension, which coincides with the topological dimension.
@@ 2.3. MCS spaces
The following definition is due to Perelman [25], [26].
@@
A metrizable space $X$ is called an MCS space of dimension $n$ if every point of $X$ has a neighborhood pointed homeomorphic to an open cone over a compact MCS space of dimension $n-1$. Here we assume that the empty set is the unique compact $(-1)$-dimensional MCS space, the cone over which is a point.
@@
Every MCS space has a natural stratification. The $k$-dimensional stratum consists of those points whose conical neighborhood topologically splits into $B^{k}\times C(L)$ but not $B^{k+1}\times C(L^{\prime})$, where $L$ and $L^{\prime}$ are compact MCS spaces. It easily follows from its inductive definition that the MCS stratification is a WCS stratification mentioned in Remark 2.3. The only reason we cannot say that it is a CS stratification is the lack of the isomorphism preserving the stratification in the normal directions as in Remark 2.3(2).
@@
The proof actually shows that every Alexandrov space is a nonbranching MCS space, which is defined in the same inductive way as in Definition 2.4, but allowing only $1$-dimensional manifolds with boundary to be $1$-dimensional nonbranching MCS spaces.
In order to prove the above theorem, Perelman [25], [26] developed the Morse theory for distance maps on Alexandrov spaces, which we will review in the next section (see also [27, §2], [20]).
@@ 2.4. Canonical neighborhoods
The following canonical neighborhood theorem is proved by reverse induction on $k$, using the deformation theorem of Siebenmann for WCS sets [34, 5.4, 6.9, 6.14]. Let $X$ be an Alexandrov space of dimension $n$.
Theorem 2.8 ([26, 1.3]).
@@
In particular, the homeomorphism $h$ is constructed from (3): the first $B^{k}$-coordinate is given by $f$ and the radial function of $C(L)$ is given by $g$.
We will not state the definition of a regular map here (see [26], [20]). It is more meaningful to give a basic example: for $k+1$ (!) points $a_{i}\in X$ ($0\leq i\leq k$), the distance map
$f=(d(a_{1},\cdot),\dots,d(a_{k},\cdot)):X\to\mathbb{R}^{k},$
is regular at $p$ if the angles $\angle a_{i}pa_{j}$ are greater than $\pi/2$ for all $0\leq i\neq j\leq k$. However, the general definition is more complicated to include the function $g$ above.
@@
Any regular map is open. Moreover, the second half of (3) is a consequence of the regularity of $(f,g)$ and the fibration theorem proved simultaneously in [26], asserting that any proper regular map is a locally trivial fibration.
@@
Given a regular map $f_{0}$ at $p$ to $\mathbb{R}^{k_{0}}$ ($k_{0}\leq k$), the map $f$ can be chosen to respect $f_{0}$, that is, the first $k_{0}$-coordinates of $f$ coincide with $f_{0}$.
@@
The above $K$ is called a canonical neighborhood of $p$ associated with the map $(f,g)$. We also denote it by $K(f,g)$ to indicate $(f,g)$. If $f$ respects $f_{0}$ in the above sense, we say that $K$ respects $f_{0}$.
@@ 2.5. Extremal subsets
An extremal subset, introduced by Perelman–Petrunin [29], is (informally) a closed subset of an Alexandrov space such that the set of normal directions at each point has diameter $\leq\pi/2$. Typical examples are
  • a point at which all angles $\leq\pi/2$, called an extremal point;
  • the closures of the MCS strata of an Alexandrov space; especially the boundary, i.e., the closure of the codimension $1$ stratum;
  • the projection of the fixed point set into the quotient space of a Riemannian manifold by an isometric compact group action; more generally, the projections of the fixed point sets of closed subgroups.
See also [27, §3], [31, §4], [11], [12] for the formal definition and properties.
We first recall how to define the extremal stratification of an Alexandrov space. The following basic properties were proved in [29].
@@
An extremal subset is called primitive if it cannot be represented as a union of two proper extremal subsets. For a primitive extremal subset $E$, its main part $\mathring{E}$ is the relative complement of all proper extremal subsets in $E$.
@@
In other words, $E$ is a minimal extremal subset containing some point $x$. The main part of $E$ is the set of those $x$ for which $E$ is the minimal extremal subset containing it.
Clearly, the main parts of all primitive extremal subsets define a disjoint covering of an Alexandrov space. Furthermore, Perelman–Petrunin [29, §2] developed the Morse theory on extremal subsets by extending the canonical neighborhood theorem 2.8 to the restriction (!) of a regular map to extremal subsets. In particular, they proved in [29, 3.8, 1.2] that
  • any primitive extremal subset is an MCS space;
  • the closures of its MCS strata are extremal subsets.
These imply that the main part of a primitive extremal subset is a (connected) topological manifold contained in some MCS stratum of the ambient Alexandrov space. Let us summarize the above as follows.
@@
The dimension of the main part of a primitive extremal subset $E$ coincides with the Hausdorff and topological dimensions of $E$, which we simply call the dimension of $E$ ([12]). It also coincides with the maximal integer $k$ such that there exists a regular map to $\mathbb{R}^{k}$ at some point of $E$ ([29, §2], [20, p.133]).
Next we discuss the relationship between extremal subsets and canonical neighborhoods. The following characterization of extremal subsets is due to Perelman [28, 2.3] (see also [13, 5.1]). Let $X$ be an Alexandrov space.
@@
The converse is true, that is, if an extremal subset $E$ intersects $K(f,g)$, then $E$ contains $g^{-1}(0)$. This follows from the openness of a regular map restricted to an extremal subset, [29, §2].
The next relative version of the canonical neighborhood theorem 2.8 with respect to extremal subsets was proved by Kapovitch [20, §9].
Complement 2.19 (to Theorem 2.8).
The homeomorphism $h$ of Theorem 2.8 can be chosen to respect (all) extremal subsets. More precisely, for any extremal subset $E$ intersecting $K$, the restriction of $h$ gives a homeomorphism to the subcone:
@@
As in Remark 2.9, this also follows from the relative fibration theorem proved simultaneously in [20, §9], asserting that any proper regular map is a locally trivial fibration respecting extremal subsets. Indeed, applying it to the regular map $(f,g)$ on $K\setminus g^{-1}(0)$ and gluing the homeomorphism $f$ on $g^{-1}(0)\subset E$ yield the desired homeomorphism $h$. The proof relies on the relative version of the deformation theorem of Siebenmann [34, 5.10, 6.10, 6.16].
@@
In the above statement, it is possible that $L\cap E=\emptyset$; in this case $K\cap E=g^{-1}(0)$ corresponds to $B^{k}\times\{o\}$, where $o$ is the vertex of the cone.
@@
Kapovitch only discussed the case of one extremal subset, but in view of the relative deformation theorem of Siebenmann cited above, it is easy to generalize it to the multiple case.
The key difference between the absolute version of the Morse theory on extremal subsets by Perelman–Petrunin [29, §2] and the relative version by Kapovitch [20, §9] is whether it includes information on the normal directions or not. The latter is essential for proving the main theorem 1.1.
We conclude this section with a well-known example from geometric topology. From our point of view, this gives an example for which the extremal stratification does not coincide with the MCS stratification.

### v2

@@ Abstract.
We prove that the extremal stratification of an Alexandrov space introduced by Perelman–Petrunin is a CS stratification in the sense of Siebenmann. We also show that every space of directions of an Alexandrov space without proper extremal subsets is homeomorphic to a sphere. In the polyhedral case, we obtain a PL structure.
@@
Alexandrov spaces, stratification, CS sets, extremal subsets, polyhedral spaces, PL manifolds, sphere theorem
@@ 1.1. Main result
The extremal stratification of an Alexandrov space introduced by Perelman–Petrunin is a CS stratification in the sense of Siebenmann.
@@
From a geometrical point of view, the proof of Theorem 1.1 is merely a combination of known technical results and provides nothing new. Nevertheless, the author believes it is worth publishing because it connects two concepts in different fields, yet even a statement cannot be found in the literature. Another motivation comes from the collapsing theory of Riemannian manifolds, where the extremal stratification of limit Alexandrov spaces plays an essential role ([32], [12], [13]; especially in the sheaf-theoretic conjectures by Alesker [1]).
@@
At the same time as [30], Perelman–Petrunin [33] introduced the notion of an extremal subset — a singular set of an Alexandrov space defined by a purely geometrical condition. By extending the Morse theory to extremal subsets, they proved that any Alexandrov space is, in a certain sense, uniquely stratified into its own extremal subsets. Here we refer to it as the extremal stratification. The extremal stratification is finer than the MCS stratification (Arrow (2)): the former reflects not only the topological singularity but also the geometrical singularity of an Alexandrov space. For example, the extremal stratification of a squire consists of the four vertices, the interiors of the four edges, and the interior of the squire itself, whereas the MCS stratification consists only of the boundary and the interior of the squire.
@@
The above-mentioned Morse theory relies heavily on the work of Siebenmann [39] on the deformation of homeomorphisms on stratified sets. In [39], Siebenmann introduced two notions of stratified sets: one is a CS set (= cone-like stratified) and the other is a WCS set (= weakly cone-like stratified). As the names suggest, CS sets are WCS sets (Arrow (3)). From our point of view, the main difference of these two is whether the local normal trivialization preserves the stratification or not (see Remark 2.3). Still, the fundamental deformation theorem of Siebenmann works in the weaker setting of WCS sets.
@@
In [29, p.6], [30, p.207], Perelman only stated that the MCS stratification of an Alexandrov space is a WCS stratification (Arrow (4)), which was enough for him to develop the Morse theory with the help of the deformation theorem of Siebenmann for WCS sets. However, there was no mention of the CS structure in [29], [30], even in [33], and as already remarked, since then there have been no papers about it (except for a closely related result of Kapovitch [23, §9] mentioned below). This is why we consider this problem, and the answer is, as stated in Theorem 1.1, given by the extremal stratification of Perelman–Petrunin (Arrow (1)).
@@
In fact, the proof of Theorem 1.1 is almost the same as the original proof of the existence of the extremal stratification by Perelman–Petrunin [33, 3.8, 1.2]. The only difference is that we make use of the relative fibration theorem with respect to extremal subsets developed later by Kapovitch [23, §9], instead of the absolute one given in [33, §2].
1.2. Further results
Now we discuss further regularity results on the extremal stratification of Alexandrov spaces.
@@ 1.2. Further results
The extremal stratification theory implies that if an Alexandrov space contains no proper extremal subsets, then it must be a topological manifold. However, even if an Alexandrov space is a topological manifold, it does not necessarily mean that each link (i.e., the space of directions) is a topological sphere. See Example 2.26 based on the double suspension theorem of Cannon [7] and Edwards [9], [10]. The infinitesimal characterization of Alexandrov topological manifolds by Wu [40] (cf. [41]) only shows that the link is a homology manifold with the homotopy type of a sphere. Regarding this problem, we prove the following.
@@
We prove the following more rigid result in Theorem 4.1: the set of points with nonspherical links in an Alexandrov topological manifold forms an extremal subset. As an application, we obtain a new sphere theorem for the link of an Alexandrov homology manifold with positive curvature (Corollary 4.6). This provides a sufficient, and in some sense optimal, condition for ruling out the double suspension over a nonsimply-connected homology sphere.
Theorem 1.2 naturally leads us to the conjecture that any Alexandrov space without proper extremal subsets would admit some kind of PL structure. Let us formulate this conjecture as follows.
Conjecture 1.3.
Suppose an Alexandrov space contains no proper extremal subsets. Then every iterated space of directions is homeomorphic to a sphere.
Here an iterated space of directions refers to a space of directions of a space of directions of … of a space of directions. The conclusion of Conjecture 1.3 holds for any noncollapsing limit of Riemannian manifolds with sectional curvature bounded below ([22, 1.4]), while the limit space may have proper extremal subsets. Another sufficient condition for Conjecture 1.3 is that every space of directions has radius greater than $\pi/2$ ([34, §3], cf. [15]). This assumption means that every direction has a $\pi/2$-apart direction, which prevents the existence of proper extremal subsets ([33, 1.4.1]). Compare with the corresponding results [26, 1.3, 1.5] in the context of geodesically complete spaces with curvature bounded above.
We shall verify Conjecture 1.3 in the case of polyhedral spaces. In this case, the conclusion of Conjecture 1.3 is equivalent to the condition that the associated triangulation is PL. See Section 5 for more details.
Theorem 1.4.
If a polyhedral Alexandrov space contains no proper extremal subsets, then the triangulation is PL.
Remark 1.5.
To the best of the author’s knowledge, it is an open problem whether an Alexandrov space admits a triangulation, even when there are no proper extremal subsets. Note that there exists a topological manifold not admitting any triangulation (see [27]).
Organization. In Section 2, we recall background material, including the definitions of the four stratifications mentioned in the introduction. In Sections 3, 4, and 5, we prove Theorems 1.1, 1.2, and 1.4, respectively. Although the proof of Theorem 1.1 is based on several technical results, all we have to do is simply combine them. For this reason, Sections 2 and 3 are rather expository. On the other hand, Section 4 is technical and demands more familiarity with Alexandrov geometry. Section 5, dealing with polyhedral spaces, requires little knowledge of Alexandrov geometry. In Appendix A, we give an example of a primitive extremal subset of codimension $2$ that is not an Alexandrov space with respect to the induced intrinsic metric. As with the main theorem, no such examples can be found in the literature.
Acknowledgments. I would like to thank Semyon Alesker for his original question regarding Theorem 1.1 that motivated me to write this paper. I am also grateful to John Harvey, Mohammad Alattar, and Shijie Gu for their valuable comments on an earlier version of this paper.
@@ 2. Preliminaries
Here we begin with the definition of a CS set and then recall several notions and results from Alexandrov geometry that will be needed to prove Theorem 1.1. In fact, once those results are assumed, the proof requires almost no geometrical knowledge.
@@ 2.1. CS sets
First we recall the definition of a CS set introduced by Siebenmann [39, §1]. For simplicity, we restrict our attention to the finite-dimensional case.
@@
An isomorphism between (not necessarily TOP) stratified sets $X$ and $Y$ is a homeomorphism between $X$ and $Y$ preserving their strata of the same dimension. This will be denoted by $X\cong Y$.
For stratified sets $X$ and $Y$, the direct product $X\times Y$ has a natural stratification. The open cone
$C(X):=X\times[0,1)/X\times\{0\}$
has a natural stratification as well, which is induced from the product structure of $X\times(0,1)$ together with the zero-dimensional stratum consisting of the vertex of the cone. An open subset $U$ of $X$ also has a natural stratification by restriction.
@@
A TOP stratified set $X$ is called a CS set if, for any $x\in X_{k}$, there exist an open neighborhood $U$ of $x$ in $X$ and a compact stratified set $L$ together with an isomorphism
@@
There is a weaker notion of a WCS set introduced by Siebenmann [39, §5]. There are two relaxations of the above conditions:
@@
Any Alexandrov space is known to be a WCS set, but not known to be a CS set in the literature. As we will see in Section 2.3, the problem lies in (2) rather than (1) (i.e., $C(L)$ is a real cone but $h$ might not be an isomorphism). In particular, we do not need the definition of the mock cone here.
@@ 2.2. Alexandrov spaces
Next we recall the definition of an Alexandrov space introduced by Burago–Gromov–Perelman [6] (though we will not use it directly). See also [5], [3], [36] for the basic theory of Alexandrov spaces.
Let $X$ be a metric space. A shortest path in $X$ is an isometric embedding of a closed interval into $X$. We say that $X$ is a geodesic space if every two points can be joined by a shortest path.
Fix $\kappa\in\mathbb{R}$. The model plane of curvature $\kappa$ refers to a sphere, Euclidean plane, or a hyperbolic plane of constant curvature $\kappa$. For a geodesic triangle $\triangle pqr$ in a metric space, the comparison triangle ~$\tilde{\triangle}pqr$ means a geodesic triangle in the model plane with the same side-lengths (if it exists and is unique up to isometry when $\kappa>0$).
Definition 2.4.
A complete geodesic space $X$ is called an Alexandrov space with curvature $\geq\kappa$ if for any geodesic triangle $\triangle pqr$ in $X$ and the comparison triangle ~$\tilde{\triangle}pqr$ in the model plane of curvature $\kappa$, the natural correspondence
$\triangle pqr\to\tilde{\triangle}pqr$
is distance-nondecreasing.
In particular, this enables us to define the angle $\angle pqr$ for every pair of shortest paths $qp$ and $qr$ (see the above references). The notion of angle plays a fundamental role in the geometry of Alexandrov spaces.
Example 2.5.
The following are Alexandrov spaces with curvature $\geq\kappa$.
  • A Riemannian manifold $M$ with sectional curvature $\geq\kappa$.
  • A convex subset and a convex hypersurface in $M$ ([2]).
  • A quotient space $M/G$, where $G$ is a compact group acting on $M$ by isometries.
  • A Gromov–Hausdorff limit of Riemannian manifolds $M_{j}$ with sectional curvature $\geq\kappa$.
In this paper, we always assume the finite-dimensionality of Alexandrov spaces. Here the dimension of an Alexandrov space means its Hausdorff dimension, which coincides with the topological dimension.
@@ 2.3. MCS spaces
Here we recall another notion of stratified space introduced by Perelman [29], [30].
@@
A metrizable space $X$ is called an MCS space of dimension $n$ if, for any $x\in X$, there exist a neighborhood $U$ of $x$ in $X$ and a compact MCS space $L$ of dimension $n-1$ together with a pointed homeomorphism
$(U,x)\to(C(L),o)$
where $o$ is the vertex of the cone $C(L)$. Here we assume that the empty set is the unique compact $(-1)$-dimensional MCS space, the cone over which is a point.
@@
Every MCS space $X$ has a natural stratification. The $k$-dimensional stratum $X_{k}$ consists of those points whose conical neighborhood topologically splits into $B^{k}\times C(L)$ but not $B^{k+1}\times C(L^{\prime})$, where $L$ and $L^{\prime}$ are compact MCS spaces. It is easy to see that the MCS stratification is a WCS stratification mentioned in Remark 2.3. More precisely, if $x\in X_{k}$, then there exist an open neighborhood $U$ of $x$ in $X$ and a compact MCS space $L$ such that there is a homeomorphism
$h:U\to\ B^{k}\times C(L)$
that sends the bottom stratum $U\cap X^{k}$ to $B^{k}\times\{o\}$, where $o$ is the vertex of $C(L)$. The reason we cannot say this is a CS stratification is that it is unclear whether $h$ preserves the higher-dimensional strata, as noted in Remark 2.3.
Perelman proved the following theorem.
@@
The proof actually shows that every Alexandrov space is a nonbranching MCS space. This is defined in the same inductive way as in Definition 2.6, but allowing only $1$-dimensional manifolds with boundary to be $1$-dimensional nonbranching MCS spaces.
In order to prove the above theorem, Perelman developed the Morse theory for distance maps on Alexandrov spaces. We will review it in the next section (see also [31, §2], [23]).
@@ 2.4. Canonical neighborhoods
The following theorem is proved by reverse induction on $k$, using the deformation theorem of Siebenmann for WCS sets [39, 5.4, 6.9, 6.14]. Let $X$ be an Alexandrov space of dimension $n$.
Theorem 2.10 ([30, 1.3]).
@@
In particular, the homeomorphism $h$ is constructed from (3): the first $B^{k}$-coordinate is given by $f$ and the projection to the second factor of $C(L)=L\times[0,1)/L\times\{0\}$ is given by $g$.
$p$$L$$f$$g$
We will not state the definition of a regular map here (see [30], [23]). It would be more helpful to give a basic example: for $k+1$ (!) points $a_{i}\in X$ ($0\leq i\leq k$), the distance map
$f=(d(a_{1},\cdot),\dots,d(a_{k},\cdot))$
is regular at $p$ if the angles $\angle a_{i}pa_{j}$ are all greater than $\pi/2$, where $0\leq i\neq j\leq k$. This condition should be regarded as linear independence of the gradients of the coordinate functions of $f$. The precise definition is more complicated to include the function $g$ in Theorem 2.10.
@@
Any regular map is open. Moreover, the second half of Theorem 2.10(3) is a consequence of the regularity of $(f,g)$ and the fibration theorem proved simultaneously in [30], asserting that any proper regular map is a locally trivial fibration.
@@
Given a regular map $f_{0}:X\to\mathbb{R}^{k_{0}}$ at $p$, the map $f$ of Theorem 2.10 can be chosen to respect $f_{0}$, that is, the first $k_{0}$-coordinates of $f$ coincide with $f_{0}$, where $k_{0}\leq k$.
Theorem 2.10 allows us to introduce the following notion, which plays a key role in the proof of Theorem 1.1.
@@
The subset $K$ of Theorem 2.10 is called a canonical neighborhood of $p$ associated with the map $(f,g)$. We also denote it by $K(f,g)$ to indicate $(f,g)$. If $f$ respects $f_{0}$ in the above sense, we say that $K$ respects $f_{0}$.
@@ 2.5. Extremal subsets
We now recall the definition of an extremal subset introduced by Perelman–Petrunin [33]. See also [31, §3], [35, §4], [36, §9.7], [11], [14] for the basic theory.
Roughly speaking, an extremal subset is a closed subset of an Alexandrov space such that the set of normal directions at each point has diameter $\leq\pi/2$. The precise definition is as follows (but we will not directly use it in this paper).
Definition 2.14.
A closed subset $E$ of an Alexandrov space $X$ is called extremal if the following holds: for any $q\in X\setminus E$, if the distance function $d(q,\cdot)$ restricted to $E$ attains a local minimum at $p\in E$, then $p$ is a critical point of $d(q,\cdot)$, i.e.,
$\angle qpx\leq\pi/2$
for every $x\in X$. By convention, $\emptyset$ and $X$ are extremal subsets.
Note that the above definition of a critical point is compatible with the regularity of a distance function in the previous section.
Example 2.15.
The following are extremal subsets.
  • A point at which all angles $\leq\pi/2$. This called an extremal point.
  • The closures of the MCS strata of an Alexandrov space (see Section 2.3); especially the boundary, i.e., the closure of the codimension $1$ stratum.
  • If a compact group $G$ acts isometrically on an Alexandrov space $X$, then $X/G$ is an Alexandrov space. The projection of the fixed point set of any closed subgroup $H$ of $G$ is an extremal subset of $X/G$.
We will recall how to define the extremal stratification of an Alexandrov space. The following basic properties were proved in [33].
@@
An extremal subset $E$ is called primitive if it cannot be represented as a union of two proper extremal subsets, i.e., if
$E=E_{1}\cup E_{2},$
where $E_{1}$ and $E_{2}$ are extremal, then $E=E_{1}$ or $E=E_{2}$. For a primitive extremal subset $E$, its main part $\mathring{E}$ is $E$ with all its proper extremal subsets removed.
@@
In other words, a primitive $E$ is a minimal extremal subset containing some point $x\in X$. The main part of $E$ is the set of those $x$ for which $E$ is the minimal extremal subset containing it.
Clearly, the main parts of all primitive extremal subsets form a disjoint covering of an Alexandrov space. Furthermore, Perelman–Petrunin [33, §2] developed the Morse theory on extremal subsets by extending Theorem 2.10 to the restriction (!) of a regular map to extremal subsets. In particular, they proved in [33, 3.8, 1.2] the following:
  • Any primitive extremal subset $E$ is an MCS space.
  • The closures of the MCS strata $E_{k}$ of $E$ are extremal subsets.
(Technically Perelman–Petrunin introduced a variant of MCS space and showed that every extremal subset is this variant, but for primitive ones, we do not need this notion.) The above two facts imply that the main part of a primitive extremal subset is a connected topological manifold and is contained in some MCS stratum of the ambient Alexandrov space. Let us summarize this as follows.
@@
The dimension of the main part of a primitive extremal subset $E$ coincides with the Hausdorff and topological dimensions of $E$. We simply call it the dimension of $E$ ([11]). It also coincides with the maximal integer $k$ such that there exists a regular map to $\mathbb{R}^{k}$ at some point of $E$ ([33, §2], [23, p.133]).
Next we discuss the relationship between extremal subsets and canonical neighborhoods. The following characterization of extremal subsets is due to Perelman [32, 2.3] (see also [12, 5.1]). Let $X$ be an Alexandrov space.
@@
The converse is true, that is, if an extremal subset $E$ intersects $K(f,g)$, then $E$ contains $g^{-1}(0)$. This follows from the openness of a regular map restricted to an extremal subset, [33, §2].
The next relative version of Theorem 2.10 with respect to extremal subsets was proved by Kapovitch [23, §9].
Complement 2.23 (to Theorem 2.10).
The homeomorphism $h$ of Theorem 2.10 can be chosen to respect all extremal subsets, i.e., for any extremal subset $E$ intersecting $K$, the restriction of $h$ gives a homeomorphism to the subcone:
@@
As in Remark 2.11, this also follows from the relative fibration theorem proved simultaneously in [23, §9], asserting that any proper regular map is a locally trivial fibration respecting extremal subsets. Indeed, applying it to the regular map $(f,g)$ on $K\setminus g^{-1}(0)$ and gluing the homeomorphism $f$ on $g^{-1}(0)\subset E$ yield the desired homeomorphism $h$. The proof relies on the relative version of the deformation theorem of Siebenmann [39, 5.10, 6.10, 6.16].
@@
In the above statement, it is possible that $L\cap E=\emptyset$; in this case $K\cap E=g^{-1}(0)$ corresponds to $B^{k}\times\{o\}$, where $o$ is the vertex of the cone $C(L)$.
@@
Kapovitch only discussed the case of one extremal subset, but in view of the relative deformation theorem of Siebenmann cited above, it is easy to generalize it to the multiple case (as observed in [20, 4.3]).
The key difference between the absolute version of the Morse theory on extremal subsets by Perelman–Petrunin [33, §2] and the relative version by Kapovitch [23, §9] is whether it includes information on the normal directions or not. The latter is essential for proving Theorem 1.1.
2.6. Double suspension
We conclude this section with a well-known example from geometric topology. From our point of view, this gives an example for which the extremal stratification does not coincide with the MCS stratification, due to the existence of nonspherical links.
For a metric space $\Sigma$ with diameter $\leq\pi$, we denote by $S(\Sigma)$ the spherical suspension over $\Sigma$, that is,
$S(\Sigma)=\Sigma\times[0,\pi]/\sim,$
where we identify each $\Sigma\times\{0\}$ and $\Sigma\times\{\pi\}$ with a point, and equip $S(\Sigma)$ with a spherical metric regarding $\Sigma\times\{\pi/2\}$ as an equator (see [5, 3.6.3] for more details). We also denote the $k$-fold spherical suspension over $\Sigma$ by
$S^{k}(\Sigma)=\underbrace{S(S(\cdots S}_{k}(\Sigma)\cdots)).$